OpenStudy (anonymous):

this deals with the ambiguous SSA case. for each find all possible solutions and sketch the triangle in each case. A=50 degreees, a=10, b=15

OpenStudy (anonymous):

Well I can help you draw the triangle but you need to use sine Law and cosine law for the rest. \[\frac{\sin50}{10} = \frac{sinB}{15}\]

OpenStudy (anonymous):

i did this and got to the same part as you but when i solve for B my calculator says domain error

OpenStudy (wolf1728):

If side a (length = 10) is opposite the 50° angle, then it seems to me that no triangle can be created.

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

You can't evaluate \[\sin^{-1}[1.149]\]

OpenStudy (anonymous):

yes exactly. so what is the answer? sorry i am just very confused

OpenStudy (wolf1728):

jake side a is too short to "seal" the triangle. (It is to short to reach another side.) The answer is no triangle can be created.